OpenStudy (anonymous):

The equation shows one mole of ethanol fuel being burned in oxygen. Convert the energy released into its equivalent mass. C2H5OH(l) + 3 O2(g) -> 2 CO2(g) + 3 H2O (l) H = -1418 kJ/mol

OpenStudy (anonymous):

@matt101

OpenStudy (anonymous):

@Somy please i need help

OpenStudy (anonymous):

Any idea?

OpenStudy (somy):

trying to remember T_T

OpenStudy (somy):

there is no T or C given so im trying to figure how to get to mass without them

OpenStudy (matt101):

We want to go from kJ/mol to kJ/g. In this case, the -1418 kJ/mol is referring to moles of C2H5OH. All you need to do is find the mass of 1 mol of C2H5OH (call that x), and then simplify the expression: -1418 kJ/x g.

OpenStudy (somy):

oh god i misunderstood the q +_+

OpenStudy (anonymous):

Hmm im still confused

OpenStudy (anonymous):

I dont understand anything about conservation of mass lol...

OpenStudy (somy):

its formula manipulation

OpenStudy (somy):

using units

OpenStudy (matt101):

You don't need to worry about conservation of mass for this question. We're just switching the units of the enthalpy. Instead of saying energy per mole of C2H5OH, we want to find energy per mass (gram) of C2H5OH. The reaction gives us the enthalpy for 1 MOLE of C2H5OH, so we just need to find the MASS of this quantity. What's the mass (in grams) of 1 mole of C2H5OH?

OpenStudy (anonymous):

6.0684 g/mol

OpenStudy (matt101):

What's the molar mass of C2H5OH?

OpenStudy (anonymous):

46.06844 g/mol

OpenStudy (somy):

\[\sf\frac{ KJ }{ mole }\] \[\sf mole = \frac{ mass }{ Mr }\] \[\sf mass = mole \times Mr\] \[\sf \frac{ KJ }{ mole } = \frac{ KJ }{ mole \times Mr }\] so to change \[\sf \frac{ KJ }{ mole }\] to \[\sf \frac{ KJ }{ g }\] you need to multiply value of \[\sf \frac{ KJ }{ mole }\] to \[\sf \frac{ 1 }{ Mr ~ of ~your~ compound }\]

OpenStudy (anonymous):

How do I do that?

OpenStudy (matt101):

Right - that's the molar mass of C2H5OH. So what's -1418 kJ/46 g?